Find the roots of the following quadratic equations by factorisation:  x ^2 – 3x – 10 = 0 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear student

Given,
x^2 – 3x – 10 =0
Taking LHS,
=>x^2 - 5x + 2x – 10
=>x(x - 5) + 2(x - 5)
=>(x - 5)(x + 2)
The roots of this equation, x^2 – 3x – 10 = 0
are the values of x for which (x - 5)(x + 2) = 0
Therefore, x - 5 = 0
or x + 2 = 0
=> x = 5 or x = -2

Thanks